# Vi-TacMan: Articulated Object Manipulation via Vision and Touch

### v1

@@ Abstract
Autonomous manipulation of articulated objects remains a fundamental challenge for robots in human environments. Vision-based methods can infer hidden kinematics but can yield imprecise estimates on unfamiliar objects. Tactile approaches achieve robust control through contact feedback but require accurate initialization. This suggests a natural synergy: vision for global guidance, touch for local precision. Yet no framework systematically exploits this complementarity for generalized articulated manipulation. Here we present Vi-TacMan, which uses vision to propose grasps and coarse directions that seed a tactile controller for precise execution. By incorporating surface normals as geometric priors and modeling directions via von Mises-Fisher distributions, our approach achieves significant gains over baselines (all $\boldsymbol{p}$<0.0001). Critically, manipulation succeeds without explicit kinematic models—the tactile controller refines coarse visual estimates through real-time contact regulation. Tests on more than $50,000$ simulated and diverse real-world objects confirm robust cross-category generalization. This work establishes that coarse visual cues suffice for reliable manipulation when coupled with tactile feedback, offering a scalable paradigm for autonomous systems in unstructured environments.
@@ I Introduction
We present Vi-TacMan, a systematic framework exploiting this complementarity. Vision detects movable and holdable parts, proposes grasps, and estimates coarse interaction directions; tactile feedback then refines execution through real-time contact regulation (Fig.˜1). Three key technical components enable robust generalization to unseen objects. First, we incorporate surface normals as geometric priors for direction estimation, providing physical constraints that significantly improve performance (pp<0.0001). Second, recognizing that multiple plausible directions may exist for unfamiliar objects, we model directional uncertainty via von Mises-Fishers on the unit sphere [16], enabling principled inference under ambiguity. Third, our detector achieves 0.86 mAP [17], reliably identifying interaction regions even in complex multi-part objects. Together, these components provide sufficient initialization required by tactile control.
@@ II-A Background: Contact-Regulating Methods
Figure 2: Inputs to the vision module of Vi-TacMan. The vision module of Vi-TacMan processes RGB-D data from a depth sensor, surface normals computed from the depth map (visualized as a normal map), and instance-level semantic masks identifying holdable and movable parts. This representation accommodates objects with multiple interactable components. Note: Holdable masks are subsets of their associated movable masks; regions appear overlapped in the visualization.
@@ III-A Dataset Preparation
To evaluate performance beyond simulation, we collect four real-world examples, each captured from five viewpoints using a Femto Bolt depth sensor. One view is illustrated in Fig.˜4(a); additional results appear in the supplementary video. These examples capture real-world diversity, including objects with single and multiple movable parts, and are reserved strictly for testing [20]. To improve depth quality, we first estimate a relative depth map using a depth foundation model [20]. Since this estimate lacks an absolute scale, we recover the correct scale by fitting a linear model between estimated disparities and ground-truth sensor measurements using RANSAC for robustness. The enhancement is illustrated in Fig.˜5.
@@ III-B Movable and Holdable Part Segmentation
Using the prepared data from Sec.˜III-A, we derive the movable and holdable masks defined in Eq.˜2, which serve as key inputs to our vision module. We train an object detector with a DINOv3 backbone and transformer-based head to detect movable and holdable parts [21, 22]. The model is trained using the AdamW optimizer with batch size 2 and learning rate $6\times 10^{-6}$. Following the protocol suggested by Lin et al[17], we report mean Average Precision across IoU thresholds from 0.50 to 0.95 (mAP@[0.50:0.95]). The model attains 0.86 mAP on the test set; detailed breakdowns appear in Tab.˜I. Since mAP above 0.6 in multi-class settings is typically considered practically useful [17] and detection is not our primary contribution, we provide the model and checkpoints in code rather than extensive baseline comparisons.
@@
Detector outputs are passed to SAM2 [23] to produce final movable and holdable masks. We associate each holdable part with its corresponding movable part by selecting the pair whose mask intersection has the largest area. Real-world results are presented in Fig.˜4(b)–(c) for illustration.
@@ III-C Grasp Selection
With movable and holdable masks defined, we establish a stable grasp on the handle. Recent advances demonstrate the effectiveness of parallel grippers for object grasping, even in cluttered environments [24, 25, 26]. The handle-grasping problem is largely simplified in our setting. We adopt a sampling-based method similar to Ten et al[27], restricting the grasp region to the holdable area. The grasping point $\boldsymbol{g}$ is defined as the centroid of this region, which determines the gripper translation. We sample gripper rotations to identify one yielding a collision-free grasp with minimal gripper width. Considering the symmetry of the parallel gripper, we select the pose closest to the robot’s home position [28]. Qualitative examples appear in Fig.˜4(d).
@@ III-D Vision-Based Displacement Estimation
where $n$ is the number of points and $\hat{\boldsymbol{q}}_{i}$ is the network’s estimate. The first term penalizes magnitude error using relative $\ell_{1}$ loss, which stabilizes optimization across a wide dynamic range and drives equal absolute errors toward zero regardless of scale. This is important because small displacements arise both outside masks and within masked regions near rotation axes. The second term aligns predicted and target directions via cosine similarity, ensuring accurate orientation even when magnitudes are small. The model is trained using AdamW optimizer with batch size 32 and learning rate $1\times 10^{-3}$.
@@ III-E Tactile Manipulation Policy
Guided by the estimated grasp and interaction direction, we deploy a tactile controller to manipulate articulated objects, following Zhao et al[14]. Due to space constraints, we refer readers to the original work for algorithmic details. A GelSight-style tactile sensor provides contact feedback [30].
@@ IV-A Simulation Studies
We evaluate interaction-direction estimation using $5,836$ test samples from categories unseen during training, as introduced in Sec.˜III-A. This setup allows us to assess generalization to previously unknown articulated objects. We compare Vi-TacMan, which leverages surface normals as an important inductive bias, against three baselines:
@@
Figure 9: Experimental platform for real-world validation.
Quantitative results are shown in Fig.˜7, where prediction error is measured as the angle between predicted and ground-truth directions. All four methods achieve median errors around $10\text{\,}\mathrm{\SIUnitSymbolDegree}$, highlighting the challenge of recovering precise motion directions on unfamiliar geometries. FlowBot3D, without modeling the distribution of point displacements, shows greater sensitivity to unseen categories. The Normal-only baseline, despite its simplicity, achieves competitive performance. Vi-TacMan reduces uncertainty by modeling the distribution of fitted normals, and explicitly incorporating surface normals further improves performance. One-sided paired t-tests confirm statistically significant improvements over all three baselines ($p<0.0001$).
@@ IV-B Real-World Experiments
As shown in Fig.˜10, Vi-TacMan guides the robot to reliably establish valid grasps on real objects and follow the estimated interaction direction. By leveraging tactile feedback, the system adapts its motions in real time, achieving consistent and robust manipulation of articulated objects. The complete manipulation process and additional experimental results are provided in the supplementary materials.
V Conclusion
We introduced Vi-TacMan, a framework for articulated object manipulation that leverages vision and touch complementarily. Rather than inferring precise but unreliable kinematics from vision alone, Vi-TacMan uses vision for coarse cues—grasp proposals and interaction direction estimates—while tactile feedback ensures robust execution. By incorporating surface normals as a geometric prior and modeling interaction directions with a vMF distribution, Vi-TacMan generalizes to unseen objects and outperforms existing baselines. Evaluations demonstrate that Vi-TacMan enables autonomous manipulation of diverse articulated objects without explicit kinematic models, highlighting the value of integrating visual guidance with tactile regulation.

### v2

@@ Abstract
Autonomous manipulation of articulated objects remains a fundamental challenge for robots in human environments. Vision-based methods can infer hidden kinematics but can yield imprecise estimates on unfamiliar objects. Tactile approaches achieve robust control through contact feedback but require accurate initialization. This suggests a natural synergy: vision for global guidance, touch for local precision. Yet no framework systematically exploits this complementarity for generalized articulated manipulation. Here we present Vi-TacMan, which uses vision to propose grasps and coarse directions that seed a tactile controller for precise execution. By incorporating surface normals as geometric priors and modeling directions via von Mises-Fisher (vMF) distributions, our approach achieves significant gains over baselines (all $\boldsymbol{p}$<0.0001). Critically, manipulation succeeds without explicit kinematic models—the tactile controller refines coarse visual estimates through real-time contact regulation. Tests on more than $50,000$ simulated and diverse real-world objects confirm robust cross-category generalization. This work establishes that coarse visual cues suffice for reliable manipulation when coupled with tactile feedback, offering a scalable paradigm for autonomous systems in unstructured environments.
@@ I Introduction
We present Vi-TacMan, a systematic framework exploiting this complementarity. Vision detects movable and holdable parts, proposes grasps, and estimates coarse interaction directions; tactile feedback then refines execution through real-time contact regulation (Fig.˜1). Three key technical components enable robust generalization to unseen objects. First, we incorporate surface normals as geometric priors for direction estimation, providing physical constraints that significantly improve performance (pp<0.0001). Second, recognizing that multiple plausible directions may exist for unfamiliar objects, we model directional uncertainty via von Mises-Fishers (vMFs) on the unit sphere [16], enabling principled inference under ambiguity. Third, our detector achieves 0.86 mAP [17], reliably identifying interaction regions even in complex multi-part objects. Together, these components provide the sufficient initialization required by tactile control.
@@ II-A Background: Contact-Regulating Methods
Figure 2: Inputs to the vision module of Vi-TacMan. The vision module of Vi-TacMan processes RGB-D data from an RGB-D sensor, surface normals computed from the depth map (visualized as a normal map), and instance-level semantic masks identifying holdable and movable parts. This representation accommodates objects with multiple interactable components. Note: Holdable masks are subsets of their associated movable masks; regions appear overlapped in the visualization.
@@ III-A Dataset Preparation
To evaluate performance beyond simulation, we collect four real-world examples, each captured from five viewpoints using a Femto Bolt RGB-D sensor. One view is illustrated in Fig.˜4(a); additional results appear in the supplementary video. These examples capture real-world diversity, including objects with single and multiple movable parts, and are reserved strictly for testing [20]. To improve depth quality, we first estimate a relative depth map using a depth foundation model [20]. Since this estimate lacks an absolute scale, we recover the correct scale by fitting a linear model between estimated disparities and ground-truth sensor measurements using RANSAC for robustness. The enhancement is illustrated in Fig.˜5.
@@ III-B Movable and Holdable Part Segmentation
Using the prepared data from Sec.˜III-A, we derive the movable and holdable masks defined in Eq.˜2, which serve as key inputs to our vision module. We train an object detector with a DINOv3 backbone and transformer-based head to detect movable and holdable parts [21, 22]. The model is trained using the AdamW optimizer on a single H100 GPU with batch size 2 and learning rate $6\times 10^{-6}$. Following the protocol suggested by Lin et al. [17], we report mean Average Precision across IoU thresholds from 0.50 to 0.95 (mAP@[0.50:0.95]). The model attains 0.86 mAP on the test set; detailed breakdowns appear in Tab.˜I. Since mAP above 0.6 in multi-class settings is typically considered practically useful [17] and detection is not our primary contribution, we provide the model and checkpoints in code rather than extensive baseline comparisons.
@@
Detector outputs are passed to SAM2 [23] to produce final movable and holdable masks on an RTX 3090 GPU. We associate each holdable part with its corresponding movable part by selecting the pair whose mask intersection has the largest area. Real-world results are presented in Fig.˜4(b)–(c) for illustration.
@@ III-C Grasp Selection
With movable and holdable masks defined, we establish a stable grasp on the handle. Recent advances demonstrate the effectiveness of parallel grippers for object grasping, even in cluttered environments [24, 25, 26]. The handle-grasping problem is largely simplified in our setting. We adopt a sampling-based method similar to Ten et al. [27], restricting the grasp region to the holdable area. The grasping point $\boldsymbol{g}$ is defined as the centroid of this region, which determines the gripper translation. We sample gripper rotations to identify one yielding a collision-free grasp with minimal gripper width. Considering the symmetry of the parallel gripper, we select the pose closest to the robot’s home position [28]. Qualitative examples appear in Fig.˜4(d).
@@ III-D Vision-Based Displacement Estimation
where $n$ is the number of points and $\hat{\boldsymbol{q}}_{i}$ is the network’s estimate. The first term penalizes magnitude error using relative $\ell_{1}$ loss, which stabilizes optimization across a wide dynamic range and drives equal absolute errors toward zero regardless of scale. This is important because small displacements arise both outside masks and within masked regions near rotation axes. The second term aligns predicted and target directions via cosine similarity, ensuring accurate orientation even when magnitudes are small. The model is trained using the AdamW optimizer on a single H100 GPU with a batch size of 32 and a learning rate of $1\times 10^{-3}$. Inference is performed on an RTX 3090 GPU for all experiments.
TABLE II: Quantitative comparison of interaction direction estimation methods on the test set. Results are reported as the mean and standard deviation of angle errors (in degrees).
Flowbot3D Normal-only Ours w/o normal Ours w/ normal
$13.92$ p m 9.45 $12.66$ p m 8.25 $10.10$ p m 6.57 $8.13$ p m 6.54
@@ III-E Tactile Manipulation Policy
Following initialization by the estimated grasp and interaction direction, all subsequent manipulation is governed by a tactile-based policy. We employ a tactile controller to manipulate articulated objects, building on the work of Zhao et al. [14, 15], which utilizes a GelSight-style tactile sensor [30] to provide contact feedback. This approach extracts tactile features from the positions of activated markers—defined as those whose normal deformation exceeds a predefined threshold. By tracking marker-wise position changes, the controller computes pose updates (Eq.˜1) via a point registration algorithm operating at $50\text{\,}\mathrm{Hz}$. For complete algorithmic details, we refer readers to the original work due to space constraints.
@@ IV Experiments
Figure 9: Experimental platform for real-world validation.
@@ IV-A Simulation Studies
We evaluate interaction direction estimation using $5,836$ test samples from categories unseen during training, as introduced in Sec.˜III-A. This setup allows us to assess generalization to previously unknown articulated objects. We compare Vi-TacMan, which leverages surface normals as an important inductive bias, against three baselines:
@@
Quantitative results are shown in Figs.˜7 and II, where prediction error is measured as the angle between predicted and ground-truth directions. All four methods achieve median errors around $10\text{\,}\mathrm{\SIUnitSymbolDegree}$, highlighting the challenge of recovering precise motion directions on unfamiliar geometries. FlowBot3D, without modeling the distribution of point displacements, shows greater sensitivity to unseen categories. The Normal-only baseline, despite its simplicity, achieves competitive performance. Vi-TacMan reduces uncertainty by modeling the distribution of fitted directions, and explicitly incorporating surface normals further improves performance. One-sided paired t-tests confirm statistically significant improvements over all three baselines ($p<0.0001$).
@@ IV-B Real-World Experiments
As shown in Fig.˜10, Vi-TacMan guides the robot to reliably establish valid grasps on real objects and follow the estimated interaction direction. By leveraging tactile feedback, the system adapts its motions in real time ($50\text{\,}\mathrm{Hz}$), achieving consistent and robust manipulation across all articulated objects. The complete manipulation process and additional experimental results are provided in the supplementary materials and on our website.
V Conclusion and Future Work
We introduced Vi-TacMan, a framework for articulated object manipulation that leverages the complementary strengths of vision and touch. Rather than inferring precise but unreliable kinematics from vision alone, Vi-TacMan uses vision for coarse guidance—grasp proposals and interaction direction estimates—while relying on tactile feedback for robust execution. By incorporating surface normals as a geometric prior and modeling interaction directions with a vMF distribution, Vi-TacMan generalizes to unseen objects and outperforms existing baselines. Our evaluations demonstrate that Vi-TacMan enables autonomous manipulation of diverse articulated objects without explicit kinematic models, highlighting the value of integrating visual guidance with tactile control.
Interpretability through hierarchical design
A key advantage of Vi-TacMan’s hierarchical architecture—which separates visual intention from tactile execution—is its inherent interpretability. Unlike end-to-end policies that map pixels directly to actions, our system explicitly generates a coarse interaction direction $\boldsymbol{d}$ and grasp $G$ before contact is made. This intermediate representation serves as a communicable “intention” that could be exposed to human users in future iterations. For instance, augmented reality projections of the intended trajectory or verbal announcements (e.g., “Opening cabinet”) prior to execution would foster safer, more predictable human-robot interaction and simplify debugging for safety certification in unstructured domestic environments.
Extending to non-rigid and multi-modal scenarios
Our current formulation leverages the Kabsch algorithm under a strict rigid-body assumption to estimate displacement. However, domestic objects often exhibit compliance or multi-stage articulation (e.g., flexible handles or nested joints). Future work will explore extending our displacement estimation to handle non-rigid deformations, potentially through deformable object tracking or sequential state estimation. Additionally, while our vMF-based modeling captures directional uncertainty, objects with multiple distinct valid interaction directions (e.g., a lever that can toggle both up and down) may require multi-modal distribution modeling or conditioning on high-level user commands.
Handling grasp failures
Our method currently assumes the initial grasp remains stable throughout manipulation. To mitigate failures from grasp slippage, we plan to integrate tactile slip detection with dynamic re-grasping primitives, enabling the system to recover from execution errors and improve overall robustness.